(1 pt) A Norman window has the shape of a rectangle surmounted by a semicircle. If the perimeter of the window is 37 feet, express the area A of the window as a function of the width x (across the base) of the window.
A(x) =

Answer by Jk22(389)

(1 pt) A Norman window has the shape of a rectangle surmounted by a semicircle. If the perimeter of the window is 37 feet, express the area A of the window as a function of the width x (across the base) of the window.
let p=perimeter, x=width in feet, h=height A=area in feet^2
p=x+2*h+Pi*x=37 => h=(37-x-Pi*x)/2
A(x)=h*x+Pi*x^2=(37-x-Pi*x)*x+Pi*x^2=37x-x*x